Shift summary for day shift 7/29/99

Sasha Glazov (glazov@hep.uchicago.edu)
Thu, 29 Jul 1999 15:36:14 -0500

Shift Summary for day shift, 29/july/99

On Shift: B.Winstein, S.Glazov

Hours of beam available at all intensities: ~8 hours
Hours of beam available at nominal intensity: ~8 hours
Hours of E832 data (special runs not included): ~6 hours 50 min
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Kfrens, Etot reset, Tapes full, problems
with ktev3.

========================================================================
Runs ended this shift:
Run 14057 16.3M events - BEAM_832
14058 1.6M events - BEAM_832
Tapes brought to FCC this shift:
QKB063-66, QKK,QKS,QKC 185-194

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
E832 datataking. KTeV4 is fixed. Problems with KTeV3.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14057
163-167 7:40 HCC is inefficient (right bottom side of CsI).
At first failed to stop triggers. Worked after
killing-restarting of ktrig.tcl. Etot reload
fixes problem.


215-216 ~8:00 Pipeline errors.

9:20 Page Dave N. to inform him that He pressure
is at 75 and falling. He new about the problem
and he promised to come in 2 hours

9:25 Sgi calls back to verify/clarify that we have
problems with ktev4. They send our call to the
experts. No estimate so far when they are going
to come.

253 9:40 ktevana errors, fixed after 2 kfrends.

11:15 SGI person promises to come, he is in fermilab

11:20 Dave N. switches temporary to helium tubes 1-2
(which are at 160 psi). He want to change to full
new tubes (5-6) this evening.

11:23 Tapes are full, run stops.
We realize that beam_muon_noktev4 configuration
is not prepared, therefore
14058 11:33 start new beam832 run

11:35 SGI apears

12:20 MCR calls, they want to stop beam in 1-1.4 hour
for repair of sagging power supply

12:28 SGI claims to repair KTeV4, its power supply is
replaced. Right after he leaves the hall
12:30 DDD3 veto full error -- End of run 14058

12:35 KTeV3 is dead. Sydney and Rick are rebooting
all machines

12:40 Call to MCR asking them to do their access now since
we are not able to take data.

12:50 Beam is gone, message "Power supl. work"

13:15 Beam is back (but not Ktev)

14060 13:24 At last start new run. KTeV4 is back. Don't
do muon run until tomorrow shutdown,

13:30 "Broken detectors" display brakes. fix_disp.
(next shift: drt_fresh_start -y)

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Do dark fresh start or keep fixing display 4 manually.